# Dynamic and Adaptive Feature Generation with LLM

### v1

@@ Abstract
The representation of feature space is a crucial environment where data points get vectorized and embedded for upcoming modeling. Thus the efficacy of machine learning (ML) algorithms is closely related to the quality of feature engineering. As one of the most important techniques, feature generation transforms raw data into an optimized feature space conducive to model training and further refines the space. Despite the advancements in automated feature engineering and feature generation, current methodologies often suffer from three fundamental issues: lack of explainability, limited applicability, and inflexible strategy. These shortcomings frequently hinder and limit the deployment of ML models across varied scenarios. Our research introduces a novel approach adopting large language models (LLMs) and feature-generating prompts to address these challenges. We propose a dynamic and adaptive feature generation method that enhances the interpretability of the feature generation process. Our approach broadens the applicability across various data types and tasks and draws advantages over strategic flexibility. A broad range of experiments showcases that our approach is significantly superior to existing methods.
00footnotetext: *Corresponding Authors.
@@ 1 Introduction
The success of machine learning (ML) algorithms generally depends on three aspects: data processing, feature engineering, and modeling. Among these, feature engineering is a critical part that directly influences the performance and effectiveness of ML models. Within feature engineering, feature generation is a vital process that transforms raw features into a structured format and optimizes the feature space by creating new features from original ones. The transformation usually involves mathematical or algorithmic operations on features. This optimization can significantly enrich the feature space, with which the ML models can draw and utilize the data information more effectively and achieve better model performance.
Given the importance of feature engineering and feature generation, there has been much work on enhancing both the effectiveness and efficiency of these methods. As deep machine learning explosively develops, automated feature engineering appears and becomes mainstream due to its convenience and remarkable performance. These auto-engineering algorithms greatly reduce the need for human calculation and determination and thus reduce subjective errors and mistakes. And in the field of automated feature generation there are outstanding works such as (Khurana et al., 2018) and (Wang et al., 2022). These methods leverage advanced techniques like reinforcement learning to generate optimal feature sets that enhance downstream tasks’ performance while maintaining a degree of explainability. Despite these methods enriching the feature space and restructuring it into a more refined space structure, like other automated feature generation and broader automated feature engineering techniques, these works have not resolved three fundamental issues:1) Lack of Explainability: Most feature engineering processes that utilize deep learning algorithms are result-oriented “black boxes”. The specific operations on features and the process of restructuring feature space are usually hard for researchers to learn and follow. 2) Limited Applicability: For different data types and downstream tasks, researchers need to manually select and switch among a wide range of feature engineering methods, which demands a high level of domain knowledge and expertise of the researchers. This limited applicability can be a significant barrier across varied ML scenarios. 3) Fixed Strategy: Existing methods usually require predefined strategies for exploration and generation. Models utilizing these methods are unable to adapt or refine their strategies based on insights acquired during the learning process of the feature space. Hence, these models lack the flexibility to dynamically adjust to the evolving understanding of the data.
@@
Figure 1: Our goal is to iteratively reconstruct the feature space to create an optimal and explainable feature set that enhances performance in downstream machine learning tasks.
@@
Our Target. We target on developing an automated feature generation method that: (1) has transparent and understandable feature generation process; (2) adapts to various data types and downstream tasks and selects best suiting operations automatically; (3) continuously learns from the feature space and refine its strategies based on new insights. For target (1) we clearly document the operations applied during the feature generation and how they transform the input data into a structured feature space. By doing so, we intend to replace the "black box" nature with a more open and interpretable framework. For target (2) we turn to large language model (LLM) that can assess the characteristics of the data and the requirements of the task and dynamically select the most appropriate strategies under proper guidance. For target (3) we develop multiple expert-level LLM agents with different strategies. These agents adjust their operations and strategies in response to changes in data patterns, task requirements, and actions of other agents.
@@
Our Method. We employ a novel approach to automate and enhance the feature generation process by adopting LLM with expert-level guidance. Our method starts by inputting an original feature set and a predefined operation set into the LLM, which then creates several expert agents. Each agent generates new features with its exclusive strategy and integrates these new features with the original features to produce enriched feature sets. The newly created feature sets are applied to downstream tasks to evaluate their performance, and the performance outcomes are fed back to the agents. The agents “communicate” and share their inferences, and reflect on the impact and effectiveness of their own and other agents’ operations on the downstream tasks. In this way the agents refine and optimize their generation strategies to better capture how variations in the feature space impact downstream tasks, furthermore reconstruct the feature space more adaptively and effectively. The iteration continues with new feature sets until the optimal feature set for the task is found or a predefined number of iterations is reached.
Contributions. In summary, our main contributions include:
@@
  • We propose LFG, a novel automated feature generation methodology based on large language model, to restructure feature space of a dataset. This end-to-end structure allows for training with LLM agents on varied feature generation strategies without manual selection of feature operations.
@@
  • We explore a dynamic and adaptive generation process based on feedback from downstream tasks. This method effectively utilizes the in-context learning and reasoning capabilities of LLMs in the feature generation process and significantly enhances the applicability and effectiveness of the generated features across various machine learning scenarios without the need to create additional ML models.
@@
  • We conduct a series of experiments to validate the effectiveness and robustness of our method (LFG) across different datasets and downstream tasks. Our experiments show that our method has clear advantages over existing methods and extensive potential of promoting to a broader range of feature engineering tasks.
2 Preliminary and Related Works
@@ 2.1 Feature Generation
Feature Engineering. Feature engineering is the process of selecting, modifying or creating new features from raw data to improve the performance of machine learning models Severyn and Moschitti (2013). The goal is to reconstruct an optimal and explainable feature representation space for a certain machine learning task (Zheng and Casari, 2018; Chen et al., 2021). This process can be described as a function ϕ::italic-$\phi:\mathcal{F}\rightarrow\mathcal{F}^{{}^{\prime}}$ that transforms the original feature set $\mathcal{F}$ into a new feature set $\mathcal{F}^{{}^{\prime}}$ with certain operations. Feature generation is a decisive category in the field of feature engineering. Feature generation means generating new attributes from existing features in a dataset through various mathematical or logical transformation operations. Feature generation methods can primarily be categorized into two main classes: (1) latent representation learning based methods, such as deep factorization machines methods (Guo et al., 2017; Song et al., 2019) and deep representation learning methods (Zhong et al., 2016; Bengio et al., 2013). These methods can create complex latent feature spaces; however, their generation processes often lack transparency and are difficult to trace and explain (Schölkopf et al., 2021). (2) feature transformation-based methods, which generate new features by performing arithmetic or aggregation operations (Nargesian et al., 2017). These methods often require a large amount of manual operations and highly rely on domain knowledge to select appropriate transformations and adjust parameters.
@@
Automated Feature Generation. With the widespread adoption of large model techniques and various deep learning methods in feature engineering field, automated feature generation has significant development and application. Automated feature generation enhances the feature space by systematically creating and integrating new features that improve the model performance (Xiang et al., 2021; Wang et al., 2022; Pan et al., 2020; Wang et al., 2022). For example,  (Katz et al., 2016) developed ExploreKit, which generates a large set of candidate features by combining information in the original features.  (Khurana et al., 2016) explores the feature space using manually crafted heuristic traversal strategies, while  (Shi et al., 2018) proposed a feature optimization approach using deep learning and feature selection to enhance traffic classification performance. These methods are more efficient than traditional manual feature engineering and enable faster processing of large datasets. However, despite their efficiency and ease of operation, these automated feature generation methods often ignore the semantic aspects of data, and their “black box” operation process make the results even harder to explain.
2.2 Large Language Models (LLMs) and Tree of Thoughts
@@
LLMs Capabilities. LLMs have revolutionized numerous fields with their extraordinary capabilities of in-context learning and step-by-step reasoning (Brown et al., 2020; Achiam et al., 2023; Wei et al., 2022a, b). With these capabilities, LLMs can effectively understand and navigate complex tasks with appropriate guidance and precise prompts. Thus they can exhibit outstanding professional competence across various specialized domains (Wang et al., 2023; White et al., 2023). Recent advancements in prompt engineering and reasoning structures have further enabled these models to tackle complex data analysis tasks and multi-step feature engineering processes (Peng et al., 2023; Liu et al., 2023).
@@
LLM Prompting. LLM prompting leverages pre-train models to recall existing knowledge and generate output based on specific guideline prompts (White et al., 2023). LLM prompting interacts closely with LLMs’ capabilities of in-context learning and step-by-step reasoning (Giray, 2023; Chu et al., 2023). Prompts with context details guide the models in thinking more coherently when handling long tasks. Moreover, step-by-step reasoning prompts can guide the model through a logical sequence of thought steps by mimicking the human-like reasoning process of solving complicated problems (Yu et al., 2023; Wei et al., 2022b; Diao et al., 2023).
@@
Tree of Thoughts. In prompting methods, one category is "logical reasoning prompting," which employs structured prompts to guide LLMs through logical analysis and problem-solving tasks. Input-Output (IO) Prompting (Wang et al., 2024; Hao et al., 2023) involves wrapping the input $x$ with task-specific instructions or examples to guide the model to produce the desired output $y$. This method is straightforward but its guiding capability is limited in complex tasks. Chain-of-Thought (CoT) Prompting (Wei et al., 2022b; Besta et al., 2024) involves a sequence of intermediate, coherent language expressions $z_{1},\ldots,z_{n}$ that logically bridge the input $x$ to the output $y$. This method enhances the model’s logical capabilities for complex problem-solving but it is limited to linear, single-strategy reasoning, which may only capture some possible solutions. Tree of Thoughts (ToT) Prompting (Yao et al., 2024) extends the CoT prompting by exploring multiple reasoning paths over thoughts. ToT works as a search over a tree structure where each node represents a partial solution within the input and contextual thoughts. ToT advantages over other methods by exploring multiple reasoning paths with different strategies. Thus it performs better in deep data analysis and multi-step feature engineering.
@@ 3 Methodology
In this section, we design a novel feature generation method called LLM Feature Generation (LFG), that is highly interpretable, dynamically adaptable, and employs an end-to-end approach to enhance the performance of downstream tasks significantly. In this method, we utilize a LLM as the auto-generator of features. We assist the LLM with the ToT technique to generate detailed interference along with each step of generation so that each decision is explainable. Furthermore, we collect the performance of downstream tasks to offer feedback to the LLM generator. Our entire workflow can be divided into two main parts: (1) Automated Feature Generation with LLM Agents; (2) Feedback and Optimization.
Figure 2: The framework of LFG. First, we input the original feature set and operation set into LLM windows. Second, we guide the LLM in creating three expert agents. Each agent generates new features with operations from the operation set, and then combines new features with original features to create a new feature subset. Then, we fed the updated feature subset into a downstream task. After that, we provide the performance of the feature subset in downstream tasks as feedback to agents and iterate the process until the best feature subset is found or the maximum number of iterations is reached.
@@ 3.1 Important Definitions
Feature Set. Let $\mathcal{D}=\{\mathcal{F},\mathcal{y}\}$ be the dataset. Here, =$\mathcal{F}=\{f_{1},f_{2},\ldots,f_{n}\}$ is the feature set, where each column represents a feature $f_{i}$, and each row represents a data sample; $y$ is the corresponding target label set for the samples. Our main purpose is to reconstruct the feature space of the dataset $\mathcal{D}=\langle\mathcal{F},\mathcal{y}\rangle$.
Operation Set. The operation set $O$ consists of mathematical operations performed on existing features to generate new ones. Operations are of two types: unary and binary. Unary operations include “square”, “exp”, “log”, etc.; binary operations include “plus”, “multiply”, “divide”, etc. The mathematical definitions of specific operations are provided in the  Appendix A.2.
Optimization Objective. We aim to construct an optimal and interpretable feature space that can enhance the performance of downstream tasks. With the feature set $\mathcal{F}$ and the operation set $\mathcal{O}$, our optimization objective is to find a reconstructed feature set $\hat{\mathcal{F}}$:
$\mathcal{F}^{*}=\underset{\hat{\mathcal{F}}}{\text{argmax}}\;\mathbf{\theta}_{\mathcal{R}}(\hat{\mathcal{F}},\mathcal{y}),$ (1)
where $\mathcal{R}$ is a downstream ML task (e.g., classification, regression, ranking, detection), $\mathbf{\theta}$ is the performance indicator of $\mathcal{R}$ and $\hat{\mathcal{F}}$ is the update of feature set $\mathcal{F}$. Here, $\hat{\mathcal{F}}$ is generated by applying operations from $\mathcal{O}$ to the original feature set $\mathcal{F}$ using a certain algorithmic structure.
@@ 3.2 Automated Feature Generation with LLM Agents
In this section, we introduce LLM agents and provide details on feature generation. We utilize LLM to generate several agents for feature-generation tasks. Here, each agent acts as an automated feature-creator, which applies specific operations on a given feature set to generate new features or delete existing features according to the prompts given in each iteration. These agents can emulate expert-level logical thinking and decision-making capabilities, hence optimizing the feature set for input data of downstream tasks and enhancing tasks’ performance.
Agents. We first define an agent $\mathcal{A}_{l}$ which operates on a feature subset $\mathcal{F}_{l}=\{f_{1},\ldots,f_{n}\}$ by applying operations to the features $f_{i}$ and $f_{j}$. The $\mathcal{F}_{l}$ is the subset of $\mathcal{F}$. The agent draws operations from an operation set $\mathcal{O}$, and these operations generate new features:
$\mathcal{F}_{l}^{{}^{\prime}}=\{\mathcal{F}_{l},g_{1},\ldots,g_{k},\ldots\}.$ (2)
@@
$g_{k}=o(f_{i},f_{j})\quad\text{for each}\quad o\in\mathcal{O}.$ (3)
Here, $\mathcal{F}_{l}$ represents the initial feature subset given to $\mathcal{A}_{l}$, $\mathcal{F}_{l}^{{}^{\prime}}$ is the output feature subset encompassing the original and all new features generated by $\mathcal{A}_{l}$, and $g_{k}$ is the new feature generated by applying the operation $o$ to features $f_{i}$ and $f_{j}$. In this way we complete one generation node $s_{l}$ of $\mathcal{A}_{l}$, and a layer $Layer_{1}$ of nodes is of total $l$ generation nodes.

### v2

@@ Abstract
The representation of feature space is a crucial environment where data points get vectorized and embedded for subsequent modeling. Thus, the efficacy of machine learning (ML) algorithms is closely related to the quality of feature engineering. As one of the most important techniques, feature generation transforms raw data into an optimized feature space conducive to model training and further refines the space. Despite the advancements in automated feature engineering and feature generation, current methodologies often suffer from three fundamental issues: lack of explainability, limited applicability, and inflexible strategy. These shortcomings frequently hinder and limit the deployment of ML models across varied scenarios. Our research introduces a novel approach adopting large language models (LLMs) and feature-generating prompts to address these challenges. We propose a dynamic and adaptive feature generation method that enhances the interpretability of the feature generation process. Our approach broadens the applicability across various data types and tasks and offers advantages in terms of strategic flexibility. A broad range of experiments showcases that our approach is significantly superior to existing methods.
@@ 1 Introduction
The success of machine learning (ML) algorithms generally depends on three aspects: data processing, feature engineering, and modeling Jordan and Mitchell (2015). Among these, feature engineering is critical, directly influencing ML models’ performance and effectiveness. Within feature engineering, feature generation is a vital process of transforming raw features into a structured format and optimizing the feature space by creating new features from the original ones Nam et al. (2024). This transformation usually involves mathematical or algorithmic operations on existing features. This optimization can significantly enrich the feature space, enabling ML models to draw and utilize data information more effectively and achieve superior performance.
@@
Figure 1: Our goal is to iteratively reconstruct the feature space to create an optimal and explainable feature set that enhances performance on downstream machine learning tasks.
Given the importance of feature engineering and feature generation, extensive research has focused on enhancing both the effectiveness and efficiency of these methodsMumuni and Mumuni (2024); Gu et al. (2024). As deep learning explosively develops, automated feature engineering has emerged as a mainstream approach due to its convenience and remarkable performance. These auto-engineering algorithms greatly reduce the need for manual calculation and determination and thus reduce subjective errors and mistakes. In the field of automated feature generation, there are outstanding works such as Khurana et al. (2018); Zhang et al. (2024b); Wang et al. (2022a). These methods leverage advanced techniques like reinforcement learning to generate optimal feature sets that enhance downstream tasks’ performance while maintaining a degree of explainability. Despite these methods enriching the feature space and restructuring it into a more refined space structure, like other automated feature generation and broader automated feature engineering techniques, these works have not resolved three fundamental issues: 1) Lack of Explainability: Most feature engineering processes that utilize deep learning algorithms are result-oriented “black boxes”. The specific operations on features and the process of restructuring feature space are usually hard for researchers to understand. 2) Limited Applicability: For different data types and downstream tasks, researchers need to manually select and switch among a wide range of feature engineering methods, which demands a high level of domain knowledge and expertise of the researchers. This limited applicability can be a significant barrier across varied ML scenarios. 3) Fixed Strategy: Existing methods usually require predefined strategies for exploration and generation. Models utilizing these methods are unable to adapt or refine their strategies based on insights acquired about the feature space during the model’s learning process. Hence, these models lack the flexibility to dynamically adjust to the evolving understanding of the data.
Our Target.
@@
We aim to develop an automated feature generation method that: (1) offers a transparent and understandable feature generation process; (2) adapts to various data types and downstream tasks by selecting best-suited operations automatically; (3) continuously learns from the feature space and refines its strategies based on new insights. For target (1), we explicitly document the operations applied during the feature generation and how they transform the input data into a structured feature space. By doing so, we intend to replace the ”black box” nature with a more open and interpretable framework. For target (2), we apply a large language model (LLM) that can assess the characteristics of the data and the requirements of the task and dynamically select the most appropriate strategies under proper guidance. For target (3), we develop multiple expert-level LLM agents with different strategies. These agents adjust their operations and strategies in response to evolving in data patterns, task requirements, and actions of other agents.
Our Method.
@@
We employ a novel approach to automate and enhance the feature generation process by adopting LLMs with expert-level guidance. Our method starts by inputting an original feature set and a predefined operation set into the LLM, which then instantiates several expert agents. Each agent generates new features with its exclusive strategy and integrates these new features with the original features to produce enriched feature sets. The newly created feature sets are applied to downstream tasks to evaluate their performance, and the performance outcomes are fed back to the agents. The agents “communicate” and share their inferences, and reflect on the impact of their own and other agents’ operations on the downstream tasks. In this way, the agents refine and optimize their generation strategies to better capture how variations in the feature space impact downstream tasks and, consequently, reconstruct the feature space more adaptively and effectively. The iteration continues with new feature sets until the optimal feature set for the task is found or a predefined number of iterations is reached.
Contributions.
Our main contributions include:
@@
  • We propose LFG, a novel automated feature generation methodology based on large language models, to restructure the feature space of a dataset. This end-to-end structure enables training with LLM agents on varied feature generation strategies without manual selection of feature operations.
@@
  • We introduce a dynamic and adaptive generation process based on feedback from downstream tasks. This method effectively utilizes the in-context learning and reasoning capabilities of LLMs in the feature generation process and significantly enhances the applicability and effectiveness of the generated features across various machine learning scenarios without the need to create additional ML models.
@@
  • We conduct a series of experiments to validate the effectiveness and robustness of our method across different datasets and downstream tasks. Our results demonstrate that our method has clear advantages over existing methods and considerable potential to promote a broader range of feature engineering tasks.
2 Preliminary and Related Work
@@ 2.1 Feature Generation
Feature engineering is the process of selecting, modifying, or creating new features from raw data to improve the performance of machine learning models Severyn and Moschitti (2013). This process can be described as a function ϕ::italic-$\phi:\mathcal{F}\rightarrow\mathcal{F}^{{}^{\prime}}$ that transforms the original feature set $\mathcal{F}$ into a new feature set $\mathcal{F}^{{}^{\prime}}$ through certain operations. Feature generation is a decisive category in the field of feature engineering. Feature generation means generating new attributes from existing features in a dataset through various mathematical or logical transformation operations. Feature generation methods can primarily be categorized into two main classes: (1) latent representation learning based methods, such as deep factorization machines  Guo et al. (2017); Song et al. (2019) and deep representation learning  Zhong et al. (2016); Bengio et al. (2013). These methods can create complex latent feature spaces; however, their generation processes often lack transparency and are difficult to trace or explain Schölkopf et al. (2021); (2) feature transformation-based methods, which generate new features by arithmetic or aggregation operations Nargesian et al. (2017); Wang et al. (2022c, b). These methods often require extensive manual operations and rely heavily on domain knowledge to select transformations and adjust parameters.
Automated Feature Generation.
@@
With the widespread adoption of large model and various deep learning methods in feature engineering field, automated feature generation is experiencing significant development. Automated feature generation enhances the feature space by systematically creating and integrating new features to improve model performance Xiang et al. (2021); Wang et al. (2022a); Pan et al. (2020); Wang et al. (2022a). For example,  Katz et al. (2016) developed ExploreKit, which generates a large set of candidate features by combining information from the original features.  Khurana et al. (2016) explores the feature space using manually crafted heuristic traversal strategies, while  Shi et al. (2018) proposed a feature optimization approach using deep learning and feature selection to enhance traffic classification performance. Although these methods are more efficient than traditional manual feature engineering and enable faster processing of large datasets, they often ignore the semantic aspects of data. Furthermore, their “black box” operation process makes the results difficult to explain.
2.2 LLMs and Tree of Thoughts
LLMs’ Capabilities.
@@ 2.2 LLMs and Tree of Thoughts
LLMs have revolutionized numerous fields with their extraordinary capabilities in in-context learning and step-by-step reasoning Xie et al. (2024); Achiam et al. (2023); Zhang et al. (2024a); Wei et al. (2022). These capabilities enable LLMs to effectively understand and navigate complex tasks given appropriate guidance and precise prompts. Thus, they exhibit professional competence across various specialized domains Wang et al. (2023); Zhang et al. (2024c, 2025a); Wang et al. (2025). Recent advancements in prompt engineering and reasoning structures have further enabled these models to tackle complex data analysis tasks and multi-step feature engineering processes Peng et al. (2023); Liu et al. (2023).
LLM Prompting.
@@
LLM prompting leverages pre-trained models to recall existing knowledge and generate output based on specific guideline prompts White et al. (2023). This process interacts closely with LLMs’ capabilities for in-context learning and step-by-step reasoning Giray (2023); Chu et al. (2023). Prompts containing context details guide models to think more coherently when handling long tasks. Moreover, step-by-step reasoning prompts can guide a model through logical thought processes by mimicking human-like reasoning to solve complex problems Yu et al. (2023); Wei et al. (2022); Diao et al. (2023).
Tree of Thoughts.
@@
In prompting methods, one category is “logical reasoning prompting,” which employs structured prompts to guide LLMs through logical analysis and problem-solving tasks. Input-Output (IO) Prompting Wang et al. (2024); Hao et al. (2023) involves wrapping the input $x$ with task-specific instructions or examples to guide the model to produce the desired output $y$. This method is straightforward but its guiding capability is limited in complex tasks. Chain-of-Thought (CoT) Prompting Wei et al. (2022); Besta et al. (2024); Zhang et al. (2025c) involves a sequence of intermediate, coherent language expressions $z_{1},\ldots,z_{n}$ that logically bridge the input $x$ to the output $y$. This method enhances the model’s logical capabilities for complex problem-solving but it is limited to linear, single-strategy reasoning, capturing only a subset of the possible solutions. Tree of Thoughts (ToT) Prompting Yao et al. (2024) extends the CoT prompting by exploring multiple reasoning paths over thoughts. ToT operates as a search over a tree structure where each node represents a partial solution within the input and contextual thoughts. ToT has advantages over other methods by exploring multiple reasoning paths with different strategies Zhang et al. (2025b). Thus, it performs better in deep data analysis and multi-step feature engineering.
@@ 3 Methodology
In this section, we design a novel feature generation method called LLM Feature Generation (LFG), which is highly interpretable, dynamically adaptable, and employs an end-to-end approach to significantly enhance downstream task performance. In this method, we utilize a LLM as an auto-generator of features. We augment the LLM with the ToT technique to generate detailed inference along with each step of generation so that each decision is explainable. Furthermore, we collect the performance of downstream tasks to offer feedback to the LLM generator. Our entire workflow is divided into two main parts: (1) Automated Feature Generation with LLM Agents; (2) Feedback and Optimization.
Figure 2: The framework of LFG. First, we input the original feature set and operation set into the LLM’s context window. Second, we guide the LLM in creating three expert agents. Each agent generates new features with operations from the operation set, and then combines new features with original features to create a new feature subset. Then, we feed the updated feature subset into a downstream task. After that, we provide the performance of the feature subset in downstream tasks as feedback to agents and iterate the process until the best feature subset is found or the maximum number of iterations is reached.
@@ 3.1 Important Definitions
Feature Set.
Let $\mathcal{D}=\{\mathcal{F},\mathbf{y}\}$ be the dataset. Here, =$\mathcal{F}=\{f_{1},f_{2},\ldots,f_{n}\}$ is the feature set, where each column represents a feature $f_{i}$, and each row represents a data sample; $\mathbf{y}$ is the corresponding target label set for the samples. Our main purpose is to reconstruct the feature space of the dataset $\mathcal{D}=\{\mathcal{F},\mathbf{y}\}$.
Operation Set.
The operation set $\mathcal{O}$ consists of mathematical operations performed on existing features to generate new ones. Operations are of two types: unary and binary. Unary operations include “square”, “exp”, “log”, etc.; binary operations include “plus”, “multiply”, “divide”, etc.
Optimization Objective.
Our objective is to construct an optimal and interpretable feature space that can enhance the performance of downstream tasks. Given the feature set $\mathcal{F}$ and the operation set $\mathcal{O}$, our optimization goal is to find a reconstructed feature set $\hat{\mathcal{F}}$:
$\mathcal{F}^{*}=\underset{\hat{\mathcal{F}}}{\text{argmax}}\;\mathbf{\theta}_{\mathcal{R}}(\hat{\mathcal{F}},\mathbf{y}),$ (1)
where $\mathcal{R}$ is a downstream ML task (e.g., classification, regression, ranking, detection), $\mathbf{\theta}$ is the performance metric of $\mathcal{R}$ and $\hat{\mathcal{F}}$ is a reconstructed feature set derived from $\mathcal{F}$. Here, $\hat{\mathcal{F}}$ is generated by applying operations from $\mathcal{O}$ to the original feature set $\mathcal{F}$ using a certain algorithmic structure.
@@ 3.2 Automated Feature Generation with LLM Agents
In this section, we introduce the concept of LLM agents and detail their role in feature generation. We utilize an LLM to generate several agents for feature-generation tasks. Here, each agent acts as an automated feature creator, which applies specific operations on a given feature set to generate new features or delete existing features based on iterative prompts. These agents can emulate expert-level logical reasoning and decision-making capabilities, thus optimizing the feature set for input data of downstream tasks and enhancing their performance.
Agents.
We first define an agent $\mathcal{A}_{l}$ that operates on a feature subset $\mathcal{F}_{l}=\{f_{1},\ldots,f_{n}\}$ by applying operations to the features $f_{i}$ and $f_{j}$. The $\mathcal{F}_{l}$ is a subset of $\mathcal{F}$. The agent draws operations from an operation set $\mathcal{O}$, to generate new features:
$\mathcal{F}_{l}^{{}^{\prime}}=\mathcal{F}_{l}\cup\{g_{1},\ldots,g_{k},\ldots\}.$ (2)
@@
$g_{k}=o(f_{i},f_{j}).$ (3)
Here, $\mathcal{F}_{l}$ is the initial feature subset for $\mathcal{A}_{l}$; $\mathcal{F}_{l}^{{}^{\prime}}$ is the output feature subset with original and all new features from $\mathcal{A}_{l}$; and $g_{k}$ is the new feature from applying operation $o$ to features $f_{i},f_{j}\in\mathcal{F}_{l}$. Thus we complete one generation node $s_{l}$ of $\mathcal{A}_{l}$, and layer $Layer_{1}$ consists of $l$ total generation nodes.